A uniform sphere of weight 200N is placed on a V- shaped wedge formed by to two flat frictionless surfaces what is the magnitude of the force exert on the sphere by each Force

1
Expert's answer
2022-01-17T12:52:52-0500

for y-direction:

"2Fcos\alpha=200"


"F=100/cos\alpha"

where "2\alpha" is angle between two flat surfaces